Mary has been working as a marketing manager and earning an annual salary of RM80,000. Using her own fixed deposits of RM100,000 from a local bank which earns her 0.3% interest per month, she decides to quit her job and open her own restaurant in Johor Bahru. Every year she pays RM24,000 for equipment leasing; RM60,000 for shop rental; and her average monthly utilities bill is RM1,000. She spends RM48,000 yearly to purchase fresh raw materials for her restaurant. Mary hires 4 workers who are paid RM2,000 each month. Her business runs very well with an average monthly sales of RM33,000. Calculate the following values on a yearly basis:

 

(i)    Implicit cost and explicit cost

(ii)  Accounting profit and economic profit



1
Expert's answer
2022-01-06T03:32:40-0500

(i)Explicit costs are calculated based on the amount of the company's expenses on external

"explicit cost=24000+60000+1000\times12+48000+2000\times12=120 000"


Implicit costs are costs that take into account the cost of internal resources. In fact, these are alternative expenses.

Implicit cost 

"i=0.003\times12=0.036"

"S=P(1+ni)=100 000(1+0.036)=103600"

Implicit costs are salary and interest:

Implicit costs=80000+3600=83600


(ii)

Accounting profit is the difference between the total revenue from the sale of products (works, services) and the explicit (external) costs, i.e. the payment for the production resources of suppliers.

Accounting profit differs from economic profit by the value of implicit (internal) costs, since economic profit does not include economic costs.

"Accounting profit=33000\times12-120000=276000"

"economic profit=33000\times12-120000-83600=192400"